# MedDiff-FM: A Diffusion-based Foundation Model for Versatile Medical Image Applications

### v2

@@ Abstract
Keywords:
Diffusion Model Foundation Model Image Denoising Anomaly Detection Super-Resolution Image Synthesis.
footnotetext: Correspondence to: Shaoting Zhang (zhangshaoting@sensetime.com), Xiaofan Zhang (xiaofan.zhang@sjtu.edu.cn).
@@ 1 Introduction
  • Through rapid fine-tuning of the diffusion foundation model via ControlNet under task-specific conditions, MedDiff-FM demonstrates effective super-resolution, lesion generation, and lesion inpainting.
2 Related Work
2.1 Patch-based Diffusion Models
@@
2.2 Diffusion Applications in Natural Images
@@
2.3 Diffusion Applications for Medical Images
@@ 3.1 Overview of MedDiff-FM
3.2 Multi-level Integrated Medical Diffusion Foundation Model
As depicted in Fig. 2, the diffusion foundation model accepts multi-level medical image inputs, specifically image-level inputs and patch-level inputs, in order to accommodate the varying resolutions and spacings of CT images. Given a 3D medical image $\mathbf{x}$, we randomly apply one of three operations: resizing the image to the patch size, randomly cropping the image to twice the patch size and then resizing it to the patch size, or randomly cropping the image to the patch size. The first operation yields image-level inputs, whereas the latter two yield patch-level inputs. The patch size used in this paper is 128 × 128 × 128. The multi-level input $\mathbf{x}_{0}$ follows the data distribution $q(\mathbf{x})$. As shown in (1), for $T$ diffusion timesteps with variances $\beta_{1},...,\beta_{T}$, it produces a sequence of noisy images $\mathbf{x}_{1},...,\mathbf{x}_{T}$, where $\mathbf{x}_{T}$ follows a standard Gaussian distribution. In the denoising process, the unconditional denoising U-Net [56] progressively transforms random noises into images via (2). In contrast, our conditional denoising U-Net takes as input the noisy image $\mathbf{\hat{x}}_{t}$, the current timestep $t$, the position embedding, and additional conditions, and outputs the estimated noise $\boldsymbol{\hat{\epsilon}}$.
@@ 3.2 Multi-level Integrated Medical Diffusion Foundation Model
During the pre-training of the diffusion foundation model, we leverage two conditioning signals, the coarse region condition $\mathbf{c}_{r}$ and the fine-grained anatomy condition $\mathbf{c}_{a}$, to achieve multi-perspective control over the denoising process and enable the model to better capture the intensity differences across different organs and tissues. The region condition $\mathbf{c}_{r}$ utilizes anatomical region classes that indicate whether the anatomical region belongs to the head and neck (HaN), chest, or abdomen. The anatomy condition $\mathbf{c}_{a}$ leverages anatomical structure masks derived from TotalSegmentator [73] and thresholding methods to impose control over anatomical structures. Since the diffusion foundation model takes multi-level medical image inputs, we aim to establish explicit spatial relationships between image-level and patch-level inputs. We adopt position encoding to obtain position embedding $\mathbf{p}_{e}$ of the X, Y, and Z coordinates, which we will discuss further in Section 3.3.
@@
$\mathbb{E}_{\mathbf{x}_{0},\boldsymbol{\epsilon},t,\mathbf{p}_{e},\mathbf{c}_{r},\mathbf{c}_{a}}\left[||\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{p}_{e},\mathbf{c}_{r},\mathbf{c}_{a})||\right].$ (3)
@@
$\mathbb{E}_{\mathbf{x}_{0},\boldsymbol{\epsilon},t,\mathbf{p}_{e},\mathbf{c}_{r},\mathbf{c}_{a},\mathbf{c}_{t}}\left[||\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,\mathbf{p}_{e},\mathbf{c}_{r},\mathbf{c}_{a},\mathbf{c}_{t})||\right].$ (4)
3.3 3D CT Image Position Embedding
@@
To construct relationships between image-level and patch-level inputs, we draw insights from Patch Diffusion [70], a patch-level diffusion model that utilizes patch coordinate conditioning. Instead of solely using patch coordinates to represent the relative position of the patch to the original image, we construct multi-level position relationships, as illustrated in Fig. 3. The multi-level position relationships establish explicit connections between the global and local information in medical images. In addition, we adapt the 2D coordinate position conditioning to 3D for medical images. The coordinate positions are pixel-level and normalized to $[-1,1]$, where $(-1,-1,-1)$ denotes the upper left back corner and $(1,1,1)$ denotes the bottom right front corner. The three position coordinate channels $\mathbf{p}_{x}$, $\mathbf{p}_{y}$, and $\mathbf{p}_{z}$ represent the X, Y, and Z coordinate positions, respectively. Furthermore, we employ the position encoding function $\text{PE}(\cdot)$ [49] to better encode positional information. The process is formulated as follows, where $L$ denotes the maximum frequency.
@@
3.4 Downstream Tasks without Fine-tuning
@@
Since LDCT images are not pure Gaussian noise, we start the reverse sampling process from an intermediate step $t_{start}$. Defining $\sigma_{n}=\bar{\sigma}_{t}=\sqrt{(1-\bar{\alpha}_{t})/\bar{\alpha}_{t}}$, substituting the estimated $\sigma_{n}$, we can obtain $t_{start}$. We then derive the expression of $\mathbf{x}_{start}$ corresponding to $t_{start}$. Given $\mathbf{y}_{0}=\mathbf{x}_{0}+\sigma_{n}\mathbf{z}$ and the diffusion process $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}$, we derive the relation in (11) and thus obtain $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{0}$. Consequently, the reverse sampling process can be initialized with $\mathbf{x}_{start}=\sqrt{\bar{\alpha}_{start}}\mathbf{y}_{0}$.
@@
3.4.2 Anomaly Detection.
Anomaly detection aims at identifying irregularities or abnormalities in medical images, aiding in disease detection and diagnosis. A large quantity of data is utilized during the pre-training of MedDiff-FM. We consider that the majority of the data contain healthy organs or tissues, while a small proportion exhibit abnormalities. Consequently, MedDiff-FM learns to synthesize normal organs or tissues. Furthermore, when using diffusion models for anomaly detection, it is more important to add noise that corrupts the abnormal regions themselves during inference. To perform anomaly detection on specific anatomical regions, we utilize anatomical structure masks to focus on the regions of interest while masking out the others. Therefore, given the masked abnormal image $\mathbf{\tilde{x}}_{0}$, we add noise at a fixed time step $t$ to obtain the corresponding noisy abnormal image $\mathbf{\tilde{x}}_{t}$. We then directly predict the original image, yielding a reconstructed healthy image $\mathbf{\tilde{x}}^{\prime}_{0}$. The anomaly map is calculated as $\mathbf{Ano_{map}}=\mathbf{\tilde{x}}_{0}-\mathbf{\tilde{x}}^{\prime}_{0}$. Finally, we use a threshold to binarize the anomaly map and obtain a binary anomaly mask $\mathbf{Ano_{mask}}$.
3.5 Downstream Tasks with Fine-tuning
3.5.1 Super-Resolution.
@@ 4 Experiments and Results
Table 1: CT datasets for the development of MedDiff-FM.
@@
4.1 Experimental Setup
4.1.1 Datasets.
MedDiff-FM datasets. We collect several publicly available medical image datasets, covering diverse anatomical regions and structures, for developing MedDiff-FM. The medical image datasets and their corresponding number of cases used in our experiments are shown in Table 1. We gather a total of 5,376 CT volumes, consisting of 362 head and neck volumes, 1,040 chest volumes, 1,732 abdomen volumes, and 2,242 whole body volumes. To balance between different anatomical regions, we only adopt 1,000 chest cases from the CT-RATE [23] dataset. These datasets are randomly divided into 90% for MedDiff-FM training and 5% for validation, with the remaining 5% reserved for the evaluation of MedDiff-FM on the downstream image synthesis task.
Table 2: CT datasets for downstream tasks.
@@ 4.1 Experimental Setup
Lesion Inpainting MED-LN [57] 90
MSD-Liver [64, 1] 131
@@
Task-specific datasets. To further validate MedDiff-FM for other downstream tasks, we adopt the datasets listed in Table 2. The Mayo 2016 dataset [47], also known as the 2016 NIH-AAPM-Mayo Clinic Low-Dose CT Grand Challenge, contains 1 mm full-dose and quarter-dose CT images from 10 patients. Covering both the chest and abdomen regions, the Mayo 2016 dataset is used to evaluate the image denoising capabilities of MedDiff-FM. The Liver CT benchmark from BMAD [2], benchmarks for medical anomaly detection, is used for anomaly detection evaluation. We utilize the RPLHR-CT [78] dataset with real-paired thick-CTs (5 mm slice thickness) and thin-CTs (1 mm slice thickness) to assess super-resolution performance. Additionally, we adopt four datasets to evaluate lesion generation and lesion inpainting tasks. Specifically, MSD-Lung and MSD-Liver are Task06 and Task03 of Medical Segmentation Decathlon (MSD) [64, 1]. MED-LN and ABD-LN are acquired from [57], containing mediastinal and abdominal lymph nodes, respectively. The four datasets are each randomly divided, with 80% used for fine-tuning MedDiff-FM, specifically for ControlNet training and validation, and the remaining 20% for testing.
4.1.2 Implementation Details.
We implement all the methods using PyTorch and carry out all the experiments using NVIDIA GeForce RTX 3090 GPUs. To obtain the anatomical structures used in MedDiff-FM, we leverage the automated whole body medical image segmentation tool, TotalSegmentator v2 [73], to segment CT images. Incorporating the 117 classes segmented by TotalSegmentator, we further derive the body class using a thresholding method, resulting in 118 classes in total. For data preprocessing, we first resample the image voxel spacing to 1.0 mm × 1.0 mm × 1.0 mm. Next, for whole body CT images, we split the head and neck, chest, and abdomen regions based on the segmentation results from TotalSegmentator and crop the images corresponding to these anatomical regions. To differentiate and extract features across different anatomical regions, we apply varying window widths and levels for window truncation to adapt to each region. For the head and neck region, the window level is set to 50 and the window width to 400; for the chest region, the window level is -500 and the window width is 1800; and for the abdomen region, the window level is 60 and the window width is 360. Subsequently, the data ranges are normalized to $[-1,1]$. The diffusion timesteps $T$ is 1,000 with cosine noise schedule. The patch size used in our experiment is 128 × 128 × 128. The model channels for MedDiff-FM and ControlNet are 32 and the number of residual blocks is 1, with the spatial transformer operating at a spatial resolution of 16 × 16 × 16. For training, we utilize $L_{1}$ loss and Adam optimizer, with a learning rate of $10^{-4}$, a batch size of 1, and 4 gradient accumulation steps. The MedDiff-FM is trained for around 150 epochs, and the ControlNet is trained for around 10k steps.
Table 3: Ablation study on the effectiveness of components in MedDiff-FM. Each component is incrementally added to the patch-level DDPM baseline.
@@
4.1.3 Evaluation Metrics.
Table 4: Comparison with state-of-the-art methods on conditional CT volume synthesis.
@@
4.2 Effectiveness Evaluation of Model Components
@@
The ablation study results are summarized in Table 3. From the perspective of generative metrics, the overall results demonstrate that incorporating multi-level integration significantly improves the performance of flexible-size image synthesis. The combination of position and position embedding further enhances image synthesis quality. The results across different anatomical regions show that, although the gains from combining position and position embedding are sometimes marginal, the overall improvements are effective. From the perspective of the Dice score in the last column, the higher Dice score indicates that the synthetic images are more consistent with the distribution of real CT images, which confirms the importance of each model component. In summary, these quantitative results highlight the effectiveness of each component in MedDiff-FM.
@@
4.3 Evaluation of Downstream Tasks without Fine-tuning
@@
The quantitative comparison results are presented in Table 4. MedDiff-FM demonstrates overall superior performance in conditional CT image synthesis with flexible sizes. Although MAISI occasionally underperforms DDPM in terms of MS-SSIM, it consistently outperforms DDPM on other metrics, indicating its advantage in generating images with better feature representations. Although MedDiff-FM shows slightly inferior performance to MAISI in terms of abdominal FID and chest MMD, it achieves superior results for multi-level flexible-size image synthesis. The Dice scores in the last column are more reliable, as higher values indicate better alignment between the synthetic images and both the given conditions and the distribution of real CT images. The qualitative results of MAISI and MedDiff-FM are shown in Fig. 5, including the axial, coronal, and sagittal views. The CT volumes generated by MedDiff-FM exhibit better continuity and finer details. These results demonstrate that the multi-level integration and position embedding in MedDiff-FM successfully facilitate flexible-size image synthesis and the construction of multi-level spatial relationships.
Table 5: Image denoising performance on the Mayo 2016 dataset. We report the PSNR and SSIM improvements of different methods over their LDCT baselines.

### v3

@@ I Introduction
  • Through efficient fine-tuning of MedDiff-FM via ControlNet under task-specific conditions, MedDiff-FM demonstrates effective super-resolution, lesion generation, and lesion inpainting.
II Related Work
II-A Patch-based Diffusion Models
@@
II-B Diffusion Applications in Natural Images
@@
II-C Diffusion Applications for Medical Images
@@ III-A Overview of MedDiff-FM
III-B Multi-level Integrated Medical Diffusion Foundation Model
As depicted in Fig. 2, the diffusion foundation model accepts multi-level medical image inputs, specifically image-level inputs and patch-level inputs, in order to accommodate the varying resolutions and spacings of CT images. Given a 3D medical image $\mathbf{x}$, we randomly apply one of three operations with equal probability: resizing the image to the patch size, randomly cropping the image to twice the patch size and then resizing it to the patch size, or randomly cropping the image to the patch size. The first operation yields image-level inputs, whereas the latter two yield patch-level inputs. The patch size used in this paper is 128 × 128 × 128. The multi-level input $\mathbf{x}_{0}$ follows the data distribution $q(\mathbf{x})$. As shown in (1), for $T$ diffusion timesteps with variances $\beta_{1},...,\beta_{T}$, it produces a sequence of noisy images $\mathbf{x}_{1},...,\mathbf{x}_{T}$, where $\mathbf{x}_{T}$ follows a standard Gaussian distribution. In the denoising process, the unconditional denoising U-Net [36] progressively transforms random noises into images via (2). In contrast, our conditional denoising U-Net takes as input the noisy image $\mathbf{\hat{x}}_{t}$, the current timestep $t$, the position embedding, and additional conditions, and outputs the estimated noise $\boldsymbol{\hat{\epsilon}}$.
@@ III-B Multi-level Integrated Medical Diffusion Foundation Model
During the pre-training of the diffusion foundation model, we leverage two conditioning signals, the coarse region condition $\mathbf{c}_{r}$ and the fine-grained anatomy condition $\mathbf{c}_{a}$, to achieve multi-perspective control over the denoising process and enable the model to better capture the intensity differences across different organs and tissues. The region condition $\mathbf{c}_{r}$ utilizes anatomical region classes that indicate whether the anatomical region belongs to the head and neck (HaN), chest, or abdomen. The anatomy condition $\mathbf{c}_{a}$ leverages anatomical structure masks derived from TotalSegmentator [37] and thresholding methods to enforce anatomical control via channel-wise concatenation. Since the diffusion foundation model takes multi-level medical image inputs, we aim to establish explicit spatial relationships between image-level and patch-level inputs. We adopt position encoding to obtain position embedding $\mathbf{p}_{e}$ of the X, Y, and Z coordinates, which we will discuss further in Section III-C.
@@
[…]
@@
[…]
III-C 3D CT Image Position Embedding
@@
To construct relationships between image-level and patch-level inputs, we draw insights from Patch Diffusion [20], a patch-level diffusion model that utilizes patch coordinate conditioning. Instead of solely using patch coordinates to represent the relative position of the patch to the original image, we construct multi-level position relationships, as illustrated in Fig. 3. The multi-level position relationships establish explicit connections between the global and local information in medical images. In addition, we adapt the 2D coordinate position conditioning to 3D for medical images. The coordinate positions are pixel-level and normalized to [1,1][-1,1], where (1,1,1)(-1,-1,-1) denotes the upper left back corner and $(1,1,1)$ denotes the bottom right front corner. The three position coordinate channels $\mathbf{p}_{x}$, $\mathbf{p}_{y}$, and $\mathbf{p}_{z}$ represent the X, Y, and Z coordinate positions, respectively. Furthermore, we employ the position encoding function $\text{PE}(\cdot)$ [38] to better encode positional information, as formulated in (6), with $L$ denoting the maximum frequency. The obtained position encoding $\mathbf{p}_{e}$ is then concatenated with the input image and incorporated into the model.
@@
III-D Downstream Tasks without Fine-tuning
@@ III-D1 Image Denoising
Since LDCT images are not pure Gaussian noise, we start the reverse sampling process from an intermediate step $t_{start}$. Let $\sigma_{n}=\bar{\sigma}_{t}=\sqrt{(1-\bar{\alpha}_{t})/\bar{\alpha}_{t}}$, and by substituting the estimated value of $\sigma_{n}$, the corresponding approximate $t$ is taken as $t_{start}$. We then derive the expression of $\mathbf{x}_{start}$ corresponding to $t_{start}$. Given $\mathbf{y}_{0}=\mathbf{x}_{0}+\sigma_{n}\mathbf{z}$ and the diffusion process $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}$, we derive the relation in (11) and thus obtain $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{0}$. Consequently, the reverse sampling process can be initialized with $\mathbf{x}_{start}=\sqrt{\bar{\alpha}_{start}}\mathbf{y}_{0}$.
@@
III-D2 Anomaly Detection
Anomaly detection aims at identifying irregularities or abnormalities in medical images, aiding in disease detection and diagnosis. A large quantity of data is utilized during the pre-training of MedDiff-FM. We consider that the majority of the data contain healthy organs or tissues, and although a small fraction exhibit abnormalities in certain anatomical regions, these regions are normal in most cases. Consequently, MedDiff-FM learns to synthesize normal organs or tissues. Furthermore, when using diffusion models for anomaly detection, it is more important to add noise that corrupts the abnormal regions themselves during inference. To perform anomaly detection on specific anatomical regions, we utilize anatomical structure masks to focus on the regions of interest while masking out the others. Therefore, given the masked abnormal image $\mathbf{\tilde{x}}_{0}$, we add noise at a fixed time step $t$ to obtain the corresponding noisy abnormal image $\mathbf{\tilde{x}}_{t}$. We then directly predict the original image, yielding a reconstructed healthy image $\mathbf{\tilde{x}}^{\prime}_{0}$. The anomaly map is calculated as $\mathbf{Ano_{map}}=\mathbf{\tilde{x}}_{0}-\mathbf{\tilde{x}}^{\prime}_{0}$. Finally, we use a threshold to binarize the anomaly map and obtain a binary anomaly mask $\mathbf{Ano_{mask}}$.
III-E Downstream Tasks with Fine-tuning
III-E1 Super-resolution
@@ IV Experiments and Results
TABLE I: CT datasets for the development of MedDiff-FM.
@@
TABLE II: CT datasets for downstream tasks.
@@
& MED-LN [60] 90
Lesion Inpainting MSD-Liver [58, 59] 131
@@
TABLE III: Ablation study on the effectiveness of components in MedDiff-FM. Each component is incrementally added to the patch-level DDPM baseline.
@@
IV-A Experimental Setup
IV-A1 Datasets
MedDiff-FM Datasets. We collect several publicly available medical image datasets, covering diverse anatomical regions and structures, for developing MedDiff-FM. The medical image datasets and their corresponding number of cases used in our experiments are shown in Table I. We gather a total of 5,376 CT volumes, consisting of 362 head and neck volumes, 1,040 chest volumes, 1,732 abdomen volumes, and 2,242 whole body volumes. To balance the number of samples across different anatomical regions, we only adopt 1,000 chest cases from the CT-RATE [47] dataset. These datasets are randomly divided into 90% for MedDiff-FM training and 5% for validation, with the remaining 5% reserved for the evaluation of MedDiff-FM on the downstream image synthesis task.
Task-specific Datasets. To further validate MedDiff-FM for other downstream tasks, we adopt the datasets listed in Table II. The Mayo 2016 dataset [55], also known as the 2016 NIH-AAPM-Mayo Clinic Low-Dose CT Grand Challenge, contains 1 mm full-dose and quarter-dose CT images from 10 patients. Covering both the chest and abdomen regions, the Mayo 2016 dataset is used to evaluate the image denoising capabilities of MedDiff-FM. The Liver CT benchmark from BMAD [56], benchmarks for medical anomaly detection, is used for anomaly detection evaluation. We utilize the real-paired RPLHR-CT [57] dataset, which contains real-world thick-CTs (5 mm slice thickness) and thin-CTs (1 mm slice thickness), to assess super-resolution performance. Additionally, we adopt four datasets to evaluate lesion generation and lesion inpainting tasks. Specifically, MSD-Lung and MSD-Liver are Task06 and Task03 of Medical Segmentation Decathlon (MSD) [58, 59]. MED-LN and ABD-LN are acquired from [60], containing mediastinal and abdominal lymph nodes, respectively. The four datasets are each randomly divided, with 80% used for fine-tuning MedDiff-FM, specifically for ControlNet training and validation, and the remaining 20% for testing.
IV-A2 Implementation Details
We implement all the methods using PyTorch and carry out all the experiments using NVIDIA GeForce RTX 3090 GPUs. To obtain the anatomical structures used in MedDiff-FM, we leverage the automated whole body medical image segmentation tool, TotalSegmentator v2 [37], to segment CT images. Incorporating the 117 classes segmented by TotalSegmentator, we further derive the body class using a thresholding method, resulting in 118 classes in total. For data preprocessing, we first resample the image voxel spacing to 1.0 mm × 1.0 mm × 1.0 mm. Next, for whole body CT images, we split the head and neck, chest, and abdomen regions based on the segmentation results from TotalSegmentator and crop the images corresponding to these anatomical regions. To differentiate and extract features across different anatomical regions, we apply varying window widths and levels for window truncation to adapt to each region. For the head and neck region, the window level is set to 50 and the window width to 400; for the chest region, the window level is -500 and the window width is 1800; and for the abdomen region, the window level is 60 and the window width is 360. Subsequently, the data ranges are normalized to [1,1][-1,1]. The diffusion timesteps $T$ is 1,000 with cosine noise schedule. The patch size used in our experiment is 128 × 128 × 128. The model channels for MedDiff-FM and ControlNet are 32 and the number of residual blocks is 1, with the spatial transformer operating at a spatial resolution of 16 × 16 × 16. For training, we utilize $L_{1}$ loss and Adam optimizer, with a learning rate of $10^{-4}$, a batch size of 1, and 4 gradient accumulation steps. The MedDiff-FM is trained for around 150 epochs, and the ControlNet is trained for around 10k steps.
TABLE IV: Comparison with state-of-the-art methods on conditional CT volume synthesis.
@@ IV-A2 Implementation Details
IV-A3 Evaluation Metrics
IV-B Effectiveness Evaluation of Model Components
@@
The ablation study results are summarized in Table III. From the perspective of generative metrics, the overall results demonstrate that incorporating multi-level integration significantly improves the performance of flexible-size image synthesis. The combination of position and position embedding further enhances image synthesis quality. The results across different anatomical regions show that, although the gains from combining position and position embedding are sometimes marginal, the overall improvements are effective. From the perspective of the Dice score in the last column, the higher Dice score indicates that the synthetic images are more consistent with the distribution of real CT images, which confirms the importance of each model component. In summary, these quantitative results highlight the effectiveness of each component in MedDiff-FM.
@@
IV-C Evaluation of Downstream Tasks without Fine-tuning
@@ IV-C1 Image Synthesis
The quantitative comparison results are presented in Table IV. MedDiff-FM demonstrates overall superior performance in conditional CT image synthesis with flexible sizes. Although MAISI occasionally underperforms DDPM in terms of MS-SSIM, it consistently outperforms DDPM on other metrics, indicating its advantage in generating images with better feature representations. Although MedDiff-FM shows slightly inferior performance to MAISI in terms of abdominal FID and chest MMD, it achieves superior results for multi-level flexible-size image synthesis. The Dice scores in the last column are more reliable, as higher values indicate better alignment between the synthetic images and both the given conditions and the distribution of real CT images. The qualitative results of MAISI and MedDiff-FM are shown in Fig. 5, including the axial, coronal, and sagittal views. The CT volumes generated by MedDiff-FM exhibit better continuity and finer details. These results demonstrate that the multi-level integration and position embedding in MedDiff-FM successfully facilitate flexible-size image synthesis and the construction of multi-level spatial relationships.
TABLE V: Image denoising performance on the Mayo 2016 dataset. We report the PSNR and SSIM improvements of different methods over their LDCT baselines.
@@
IV-C2 Image Denoising
We evaluate the image denoising capabilities of MedDiff-FM on the Mayo 2016 dataset using the official 25% dose CT images. We utilize the off-the-shelf MedDiff-FM for image denoising as introduced in Section III-D1. The standard deviation of the additive noise $\sigma_{n}$ is estimated from the non-test cases and is approximately 0.15. And the regularization parameter $\lambda$ is set to 10. The results of the comparison methods in Table V are cited from [15]. Due to the different preferences for window width and level settings across methods, we report the PSNR and SSIM improvements of each method over its corresponding LDCT baseline under its respective window width and level, using the identical test cases. Specifically, the LDCT baseline under the window of [1024,3072][-1024,3072] HU yields a PSNR of 40.41 dB and an SSIM of 0.922, and the LDCT baseline under the window of [160,240][-160,240] HU achieves a PSNR of 23.10 dB and an SSIM of 0.757. MedDiff-FM performs image denoising under the window of [160,240][-160,240] HU. As shown in Table V, MedDiff-FM even significantly outperforms supervised comparison methods in terms of SSIM, while maintaining competitive PSNR performance. Unlike all other comparison methods, MedDiff-FM is applied to image denoising without additional training on the Mayo 2016 dataset, thus reducing the consumption of spatio-temporal resources. The qualitative results in Fig. 6 intuitively demonstrate that MedDiff-FM achieves effective denoising and enhances image quality, highlighting its strong capabilities in image denoising.